OpenStudy (anonymous):

find the length of the hypotenuse of a right triangle if a=1 and b= 7

OpenStudy (anonymous):

@it_mee

OpenStudy (anonymous):

Tag me if you want help.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sqrt { a^2+b^2} = \sqrt{ 1^2+7^2} =?\]

OpenStudy (anonymous):

The formula for the hypotenuse\[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

\[1^{2}=1\]\[7^{2}=49\]\[49+1=50\]\[\sqrt{50}=?\]@it_mee